# Physics-Conditioned Diffusion Models for Lattice Gauge Theory

### v1

@@ 1 Introduction
Despite the success of MCMC in simulating lattice field theories, one often encounters the issue of critical slowing down Wolff:1989wq . This phenomenon occurs when the correlation time of a system increases significantly as it approaches a critical point in parameter space, leading to escalating computational costs and insufficient diversity in the sampled configurations. Although alternative methods Wolff:1988uh ; Hasenbusch:2001ne ; Hasenbusch:2002ai ; Prokofev:2001ddj ; Brannick:2007ue ; Albergo:2019eim ; Albandea:2021lvl ; Eichhorn:2023uge have been proposed to mitigate critical slowing down in specific models, it remains a longstanding challenge in lattice simulations of Quantum Chromodynamics.
@@
Generative models in machine learning (ML), which aim to map samples from a simple prior distribution to a complex physical distribution, have shown great promise in addressing challenges such as critical slowing down in lattice field theory. Recent advancements in generative models have introduced novel approaches for lattice field theory simulations Cranmer:2023xbe ; Zhou:2023pti ; Wang:2020hji . Among these, flow-based models have emerged as a leading method for explicit likelihood estimation, offering advantages such as invertibility and gauge equivariance, which enable efficient global sampling Kanwar:2020xzo ; Boyda:2020hsi ; Kanwar:2021wzm ; Albergo:2021vyo ; Cranmer:2023xbe ; Kanwar:2024ujc ; Chen:2022ytr . Furthermore, extensions of normalizing flows, including continuous normalizing flows chen:2018 ; deHaan:2021erb ; Gerdes:2022eve ; Caselle:2023mvh and stochastic normalizing flows wu2020stochastic ; Caselle:2022acb ; Bulgarelli:2024brv , expand their capabilities in lattice simulations. However, in flow-based models, the training complexity increases significantly with larger lattice sizes DelDebbio:2021qwf ; Abbott:2022zsh ; Komijani:2023fzy ; Abbott:2023thq . Moreover, the use of the Kullback-Leibler divergence as the optimization objective can lead to mode collapse Nicoli:2023qsl ; Kanaujia:2024zrq , where the learned distribution fails to capture the full complexity of the target distribution. These challenges underscore the need to explore alternative generative models to further enhance sampling efficiency and reliability.
@@
Recently, diffusion models (DMs) have gained significant attention for their success in generating high-quality samples across diverse domains sohl-dickstein:2015deep ; ho:2020denoising ; song2019generative ; Song:2021scorebased ; yang2023diffusion ; croitoru2023diffusion , including applications in high-energy physics Mikuni:2022xry ; Amram:2023onf ; Mikuni:2023dvk ; Devlin:2023jzp . Diffusion models iteratively denoise random noise into realistic configurations through probabilistic steps bishop2023deep . Their application to lattice field theory was pioneered in Refs. Wang:2023exq ; Wang:2023sry for scalar fields, establishing the connection with stochastic quantization Parisi:1980ys ; Damgaard:1987rr ; Namiki:1993fd ; Fukushima:2024oij . Subsequent studies introduced a Feynman path integral formulation Hirono:2024zyg and analysed higher-order cumulants  Aarts:2024rsl , further solidifying the theoretical foundation of diffusion models in this context. An application to data generated using complex Langevin dynamics can be found in Ref. Habibi:2024fbn . In this work, we propose a novel framework based on diffusion models to address challenges in lattice gauge theory simulations Zhu:2024kiu . To validate our approach, we conduct experiments in two-dimensional U(1) gauge theory, demonstrating that our model effectively samples topological quantities across varying coupling constants and lattice sizes without requiring additional training.
@@
The paper is organized as follows. In Section 2 we discuss the issue of topological freezing, with subsections covering the Hybrid Monte Carlo (HMC) approach and topological charge. Section 3 introduces diffusion models and their training objectives, including the noise conditional score network and the U-Net architecture. In Section 4, we present our physics-conditioned sampler, which integrates the diffusion model with the notion of stochastic quantization and Langevin dynamics for conditional sampling while ensuring exactness. Section 5 details our numerical experiments, including data preparation, model setups, and evaluations across various lattice sizes and coupling constants. Section 6 concludes. More details are discussed in the Appendices.
@@ 2 Simulation of Lattice Gauge Fields
Abelian gauge theories in two Euclidean dimensions represent quantum electrodynamics (QED) on a plane and provide simplified models for gauge field theories Gattringer:2010zz ; Greensite:2011zz ; Kanwar:2020xzo ; Crean:2024nro , while retaining some of the essential features of nonabelian theories in four dimensions: fermion confinement Coleman:1975pw and separated topological sectors Greensite:2011zz . In the two-dimensional U(1) gauge theory, a gauge-invariant lattice action can be defined as
$S=-\beta\sum_{x}\text{Re}(U_{x,\Box}),$ (3)
@@
On an $L\times L$ square lattice with periodic boundary conditions, a gauge field on lattice site $x$ in the direction $\mu$ is represented by a link variable $U_{x,\mu}=e^{i\phi_{x,\mu}}$, where $\phi_{x,\mu}$ is the gauge angle for the link. The plaquette, or the $1\times 1$ Wilson loop ($W_{1\times 1}$), is a gauge-invariant observable. Here $\beta$ is the inverse coupling that controls the gauge interaction strength and ensures consistency with the continuum limit (at large $\beta$) in simulations Gattringer:2010zz .
@@ 2.1 Hybrid Monte Carlo
In lattice field theory, physical properties are estimated via the expectation value of observables $\hat{O}$, representing the statistical expectation over all configurations in the physics distribution. The HMC algorithm is widely used to sample configurations efficiently by introducing auxiliary momenta, and utilizing Hamiltonian dynamics Duane:1987de ; Hasenbusch:2001ne ; Hasenbusch:2002ai ; Gattringer:2010zz ; Neal:2011mrf . This enables effective phase space exploration and reduces autocorrelations, overcoming inefficiencies of random walk-based methods.
@@ 2.2 Topological Charge Freezing
In gauge theories, the topological charge plays a key role in phenomena like confinement Greensite:2011zz . It is quantized, characterizing different vacuum states, and is linked to flux quantization and instantons Greensite:2011zz . The Atiyah-Singer Index Theorem Atiyah:1968mp states that $Q$ must be an integer and remain unchanged under continuous field transformations. Thus, topological charges reside in distinct sectors separated by energy barriers Eichhorn:2023uge ; Durr:2024dal .
@@
Topological freezing refers to the phenomenon where the topological charge $Q$ remains constant over long Monte Carlo trajectories, preventing adequate sampling of different topological sectors Kanwar:2020xzo ; Albandea:2021kwe . This issue may arise as a general feature in Monte Carlo simulations of lattice theories with non-trivial topological properties Alles:1996vn ; DelDebbio:2002xa ; DelDebbio:2004xh ; Bonanno:2018xtd ; Eichhorn:2023uge . It is refereed to the critical slowing down of topological modes, where the autocorrelation time $\tau_{Q}$ grows exponentially with the correlation length $\xi_{c}$ as $\tau_{Q}\sim\exp(c\xi_{c})$. Since $\xi_{c}$ is inversely related to the square root of the string tension, $\xi_{c}\sim\sigma_{s}^{-1/2}$, it diverges as the lattice spacing $a$ decreases (or equivalently, as $\beta$ increases), approaching the continuum limit Schaefer:2010hu . Consequently, transitions between topological sectors become exponentially rare.
As illustrated in Fig. 1, see also Figs. 1 of Refs. Kanwar:2020xzo ; Albandea:2021kwe , samples within the same Markov chain (with different chains represented by different colors) remain confined to a limited number of topological sectors. This significantly hinders simulation efficiency and the accuracy of physical observables related to topology. In HMC, when the system is trapped in a restricted region of configuration space, transitions between isolated topological sectors are further suppressed Alles:1996vn ; DelDebbio:2001sj . At large $\beta$, the action landscape may develop deep, narrow wells corresponding to these separated sectors, but Markov chains, which typically evolve by making small updates to state variables, struggle to transition between them 2001Annealed ; Eichhorn:2023uge . To address this challenge, improvements to existing algorithms or alternative approaches are necessary.
@@ 3.1 Diffusion Models
Diffusion models generate samples by reversing the process of adding noise, with the goal of recovering the original distribution from pure noise bishop2023deep . While theoretically possible if the forward stochastic dynamics are known, in practice it is often difficult to determine the reverse process. To solve this problem, a deep model approximates the gradient of the logarithm of the data distribution, which feeds directly in the denoising process. The well-trained deep model on prepared lattice field configurations can serve as an effective drift term to generate more field configurations by evolving a stochastic differential equation Wang:2023exq . It can provide an efficient and global sampler to simulate lattice quantum field theories Wang:2023exq ; Wang:2023sry ; Aarts:2024rsl ; Aarts:2024agm .
@@
where the Langevin time $\xi$ goes from 00 to $T\equiv 1$, $\eta$ is the noise term, satisfying $\langle\eta(\xi)\eta(\xi^{\prime})\rangle=\delta(\xi-\xi^{\prime})$, and $f(\phi,\xi)$ is the drift term. The square of $g(\xi)$ is the time-dependent scalar diffusion coefficient. In our case, the forward drift term $f(\phi,\xi)$ is set to zero and the variance-expanding scheme is chosen, see, e.g., Ref. song2020score , with $g(\xi)=\sigma^{\xi}$. This way, the transition kernel in the forward process remains Gaussian, due to the additivity of Gaussian white noise,
@@
In this setup, we can add noise at any $0<\xi\leq T$, to achieve arbitrarily perturbed data.
The reverse denoising process corresponding to Eq. (7) is accomplished by a reverse SDE, representing an evolution backwards in time anderson:1982reversetime ,
@@
where the reverse time $t$ runs from $T$ to 00 and $\bar{\eta}$ is a noise term in the reverse time direction, still satisfying $\langle\bar{\eta}(t)\bar{\eta}(t^{\prime})\rangle=\delta(t-t^{\prime})$. Importantly, the drift term in Eq. (9) includes the gradient of the logarithm of $p_{t}(\phi)$, the probabilistic distribution at time $t$ in the forward diffusion process.
@@ 3.2 Noise Conditional Score Networks
Typically, the gradients $\nabla_{\phi}\log p_{t}(\phi)$ in Eq. (9) are unknown. However, one can model a score function, $s(\phi,\xi)$, to match the gradients song2020score . Due to demands of the scalability and ability to efficiently handle high-dimensional parameter spaces, we employ a deep neural network to model the score function. To further improve sampling efficiency, we introduce a noise scheme for the forward process, in which $\xi_{i}$ ($i=1,\ldots,N_{T}$) is a positive geometric sequence that satisfies $\xi_{1}/\xi_{2}=\cdots=\xi_{N_{T}-1}/\xi_{N_{T}}>1$.
@@
$\forall\xi\in\{\xi_{i}\}_{i=1}^{N_{T}}:\quad s_{\theta}(\phi,\xi)\approx\nabla_{\phi}\log p_{\xi}(\phi_{\xi}|\phi_{0}),$ (10)
@@
$\ell(\theta;\xi)\triangleq\frac{1}{2}\mathbb{E}_{p_{\text{data}}(\phi)}\mathbb{E}_{\tilde{\phi}\sim\mathcal{N}(\phi,\sigma_{\xi}^{2}I)}\left[\left\lVert s_{\theta}(\tilde{\phi},\xi)+\frac{\tilde{\phi}-\phi}{\sigma_{\xi}^{2}}\right\rVert_{2}^{2}\right].$ (11)
@@
where $\lambda(\xi)=\sigma_{\xi}^{2}$ song2020score . Under this choice, we have
$\lambda(\xi)\ell(\theta;\xi)=\frac{1}{2}\mathbb{E}\left[\left\lVert\sigma_{\xi}s_{\theta}(\tilde{\phi},\xi)+\frac{\tilde{\phi}-\phi}{\sigma_{\xi}}\right\rVert_{2}^{2}\right].$ (13)
Since $(\tilde{\phi}-\phi)/\sigma_{\xi}\sim\mathcal{N}(0,I)$ and $\lVert\sigma_{\xi}s_{\theta}(\phi,\xi)\rVert_{2}\propto 1$, we can deduce that the order of magnitude of $\lambda(\xi)\ell(\theta;\xi)$ is independent of $\xi$.
With regard to the specific network architecture, similar as in Ref. Wang:2023exq , we adopt a U-Net architecture as shown in Fig. 3 to model the score function. It employs an encoder-decoder structure with symmetric skip connections. The receptive field grows as the network deepens, but each convolution acts locally, making it adaptable to arbitrarily large input sizes.
Figure 3: U-Net architecture. The gray lines with arrows denote skip connections between the encoding and decoding paths. In this study, the default input shape is specified as $(\cdot,2,16,16)$, where the first dimension represents the batch size, the second corresponds to the channel dimension, and the last two denote the spatial dimensions of the field configuration.
Input and Output: The input tensor is defined as $X\in\mathbb{R}^{N\times 2\times L\times L},$ where $N$ denotes the batch size, $2$ represents the two gauge link channels, and $L$ (default $16$) is the spatial lattice size. The output tensor $Y$ has the same shape and approximates the score function, $s_{\theta}(\phi,\xi).$ For sampling, the lattice size $L$ can be adjusted accordingly.
U-Net Architecture: The encoder applies convolutional layers,
$h^{(l+1)}=\sigma\Big{(}W^{(l)}*h^{(l)}+b^{(l)}\Big{)}.$ (14)
Each layer follows group normalization, $\hat{h}^{(l)}={(h^{(l)}-\mu_{G})}/{\sqrt{\sigma_{G}^{2}+\epsilon}}.$ The decoder utilizes transposed convolutions,
@@
where $E^{(l-1)}$ indicates the corresponding encoder layer. Finally, the network output is normalized by the standard deviation of the perturbation kernel.
Locality and Periodic Padding: The lattice gauge action is constructed from Wilson loops, ensuring local interactions. Accordingly, the fully convolutional network (FCN) computes outputs locally, as
@@
It denotes the collection (or local patch) of input values that lie within a neighbourhood $\mathcal{N}(i)$ around index $i$ and $\mathcal{N}(j)$ around index $j$. Periodic boundary conditions are enforced by using periodic padding, meaning that if an index exceeds the lattice boundaries, it wraps around according to the modulo operation.

### v2

@@ 1 Introduction
Despite the success of MCMC in simulating lattice field theories, one often encounters the issue of critical slowing down Wolff (1990). This phenomenon occurs when the correlation time of a system increases significantly as it approaches a critical point in parameter space, leading to escalating computational costs and insufficient diversity in the sampled configurations. Although alternative methods Wolff (1989); Hasenbusch (2001); Hasenbusch and Jansen (2003); Prokof’ev and Svistunov (2001); Brannick et al. (2008); Albergo et al. (2019); Albandea et al. (2021); Eichhorn et al. (2024) have been proposed to mitigate critical slowing down in specific models, it remains a longstanding challenge in lattice simulations of Quantum Chromodynamics.
@@
Generative models in machine learning (ML), which aim to map samples from a simple prior distribution to a complex physical distribution, have shown great promise in addressing challenges such as critical slowing down in lattice field theory. Recent advancements in generative models have introduced novel approaches for lattice field theory simulations Cranmer et al. (2023); Zhou et al. (2024); Wang et al. (2022). Among these, flow-based models have emerged as a leading method for explicit likelihood estimation, offering advantages such as invertibility and gauge equivariance, which enable efficient global sampling Kanwar et al. (2020); Boyda et al. (2021); Kanwar (2021); Albergo et al. (2021); Cranmer et al. (2023); Kanwar (2024); Chen et al. (2023). Furthermore, extensions of normalizing flows, including continuous normalizing flows Chen et al. (2018); de Haan et al. (2021); Gerdes et al. (2023); Caselle et al. (2024) and stochastic normalizing flows Wu et al. (2020); Caselle et al. (2022); Bulgarelli et al. (2025), expand their capabilities in lattice simulations. However, in flow-based models, the training complexity increases significantly with larger lattice sizes Del Debbio et al. (2021); Abbott and others (2023a); Komijani and Marinkovic (2023); Abbott and others (2023b). Moreover, the use of the Kullback-Leibler divergence as the optimization objective can lead to mode collapse Nicoli et al. (2023); Kanaujia et al. (2024), where the learned distribution fails to capture the full complexity of the target distribution. These challenges underscore the need to explore alternative generative models to further enhance sampling efficiency and reliability.
@@
Recently, diffusion models (DMs) have gained significant attention for their success in generating high-quality samples across diverse domains Sohl-Dickstein et al. (2015); Ho et al. (2020); Song and Ermon (2019); Song et al. (2020); Yang et al. (2023); Croitoru et al. (2023), including applications in high-energy physics Mikuni and Nachman (2022); Amram and Pedro (2023); Mikuni et al. (2023); Devlin et al. (2024). Diffusion models iteratively denoise random noise into realistic configurations through probabilistic steps Bishop and Bishop (2023). Their application to lattice field theory was pioneered in Refs. Wang et al. (2024, 2023) for scalar fields, establishing the connection with stochastic quantization Parisi et al. (1981); Damgaard and Huffel (1987); Namiki (1993); Fukushima and Kamata (2025). Subsequent studies introduced a Feynman path integral formulation Hirono et al. (2024) and analysed higher-order cumulants  Aarts et al. (2025b), further solidifying the theoretical foundation of diffusion models in this context. An application to data generated using complex Langevin dynamics can be found in Ref. Habibi et al. (2025). In this work, we propose a novel framework based on diffusion models to address challenges in lattice gauge theory simulations Zhu et al. (2024). To validate our approach, we conduct experiments in two-dimensional U(1) gauge theory, demonstrating that our model effectively samples topological quantities across varying coupling constants and lattice sizes without requiring additional training.
@@
The paper is organized as follows. In Section 2 we discuss the issue of topological freezing, with subsections covering the Hybrid Monte Carlo (HMC) approach and topological charge. Section 3 introduces diffusion models and their training objectives, including the noise conditional score network and the Net architecture. In Section 4, we present our physics-conditioned sampler, which integrates the diffusion model with the notion of stochastic quantization and Langevin dynamics for conditional sampling while ensuring exactness. Section 5 details our numerical experiments, including data preparation, model setups, and evaluations across various lattice sizes and coupling constants. Section 6 concludes. More details are discussed in the Appendices.
@@ 2 Simulation of Lattice Gauge Fields
Abelian gauge theories in two Euclidean dimensions represent quantum electrodynamics (QED) on a plane and provide simplified models for gauge field theories Gattringer and Lang (2010); Greensite (2011); Kanwar et al. (2020); Crean et al. (2024), while retaining some of the essential features of nonabelian theories in four dimensions: fermion confinement Coleman et al. (1975) and separated topological sectors Greensite (2011). In the two-dimensional U(1) gauge theory, a gauge-invariant lattice Euclidean action can be defined as
$S_{E}=-\beta\sum_{x}\text{Re}(U_{x,\Box}),$ (3)
@@
On an $L\times L$ square lattice with periodic boundary conditions, a gauge field on lattice site $x$ in the direction $\mu$ is represented by a link variable $U_{x,\mu}=e^{i\phi_{x,\mu}}$, where $\phi_{x,\mu}$ is the gauge angle for the link. The plaquette, or the $1\times 1$ Wilson loop ($W_{1\times 1}$), is a gauge-invariant observable. Here $\beta$ is the inverse coupling that controls the gauge interaction strength and ensures consistency with the continuum limit (at large $\beta$) in simulations Gattringer and Lang (2010).
@@ 2.1 Hybrid Monte Carlo
In lattice field theory, physical properties are estimated via the expectation value of observables $\hat{O}$, representing the statistical expectation over all configurations in the physics distribution. The HMC algorithm is widely used to sample configurations efficiently by introducing auxiliary momenta, and utilizing Hamiltonian dynamics Duane et al. (1987); Hasenbusch (2001); Hasenbusch and Jansen (2003); Gattringer and Lang (2010); Neal (2011). This enables effective phase space exploration and reduces autocorrelations, overcoming inefficiencies of random walk-based methods.
@@ 2.2 Topological Charge Freezing
In gauge theories, the topological charge plays a key role in phenomena like confinement Greensite (2011). It is quantized, characterizing different vacuum states, and is linked to flux quantization and instantons Greensite (2011). The Atiyah-Singer Index Theorem Atiyah and Singer (1968) states that $Q$ must be an integer and remain unchanged under continuous field transformations. Thus, topological charges reside in distinct sectors separated by energy barriers Eichhorn et al. (2024); Rouenhoff and Dürr (2025).
@@
Topological freezing refers to the phenomenon where the topological charge $Q$ remains constant over long Monte Carlo trajectories, preventing adequate sampling of different topological sectors Kanwar et al. (2020); Albandea et al. (2022). This issue may arise as a general feature in Markov Chain Monte Carlo simulations of lattice theories with non-trivial topological properties Alles et al. (1996); Del Debbio et al. (2002b, 2004); Bonanno et al. (2019); Eichhorn et al. (2024). It is refereed to divergence of the autocorrelation time $\tau_{Q}$. Consequently, transitions between topological sectors become exponentially rare.
As illustrated in Fig. 1, see also Figs. 1 of Refs. Kanwar et al. (2020); Albandea et al. (2022), samples within the same Markov chain (with different chains represented by different colors) remain confined to a limited number of topological sectors. This significantly hinders simulation efficiency and the accuracy of physical observables related to topology. In traditional Markov chain Monte-Carlo methods, when the system is trapped in a restricted region of configuration space, transitions between isolated topological sectors are further suppressed Alles et al. (1996); Del Debbio et al. (2002a), situations exist even in hybrid methods like HMC, which is the most common way to sample QCD configurations. Specifically at large $\beta$, the action landscape may develop deep, narrow wells corresponding to these separated sectors, but Markov chains, which typically evolve by making small updates to state variables, struggle to transition between them Neal (2001); Eichhorn et al. (2024). To address this challenge, improvements to existing algorithms or alternative approaches are necessary.
@@ 3.1 Diffusion Models
Diffusion models generate samples by reversing the process of adding noise, with the goal of recovering the original distribution from pure noise Bishop and Bishop (2023). While theoretically possible if the forward stochastic dynamics are known, in practice it is often difficult to determine the reverse process. To solve this problem, a deep model approximates the gradient of the logarithm of the data distribution, which feeds directly in the denoising process. The well-trained deep model on prepared lattice field configurations can serve as an effective drift term to generate more field configurations by evolving a stochastic differential equation Wang et al. (2024). It can provide an efficient and global sampler to simulate lattice quantum field theories Wang et al. (2024, 2023); Aarts et al. (2025b, c). Here, a global sampler refers to a sampling method capable of proposing new configurations that differ globally from previous ones, effectively reducing the issue of long autocorrelation times and topological freezing commonly encountered in local update algorithms.
@@
where the Langevin time $\xi$ goes from 0 to $T\equiv 1$, $\eta$ is the noise term independently added to all field degrees of freedom in the field configurations $\phi$, satisfying $\langle\eta(\xi)\eta(\xi^{\prime})\rangle=\delta(\xi-\xi^{\prime})$, and $f(\phi,\xi)$ is the vector drift term. The square of the scalar $g(\xi)$ is the time-dependent scalar diffusion coefficient. In our case, the forward drift term $f(\phi,\xi)$ is set to zero and the variance-expanding scheme is chosen, see, e.g., Ref. Song et al. (2020), with $g(\xi)=\sigma^{\xi}$. This way, the transition kernel in the forward process remains Gaussian, due to the additivity of Gaussian white noise,
@@
In this setup, we can add noise at any $0<\xi\leq T$, to achieve arbitrarily perturbed data without integration. In our implementation, we choose a geometric sequence of $\xi$ to be training points.
The reverse denoising process corresponding to Eq. (7) is accomplished by a reverse SDE, representing an evolution backwards in time Anderson (1982),
@@
where the reverse time $t$ runs from $T$ to 0 and $\bar{\eta}$ is a noise term in the reverse time direction, still satisfying $\langle\bar{\eta}(t)\bar{\eta}(t^{\prime})\rangle=\delta(t-t^{\prime})$. Importantly, the drift term in Eq. (9) includes the gradient of the logarithm of $p_{t}(\phi)$, the probabilistic distribution at time $t$ in the forward diffusion process.
In Eqs. (3.1) and (3.3), the variables $\xi$ and $t$ are defined as continuous parameters in the interval $[0,1]$ to describe the underlying diffusion process. In practical implementation, we evaluate these parameters at discrete noise levels $\xi_{i}$ following a geometric sequence (e.g., $\xi_{i}=\xi_{0}\cdot r^{i}$). It is important to note that Eq. (3.2) defines the analytical marginal noise scale $\sigma_{\xi}$ for any given intensity $\xi$, rather than being derived from a path-dependent integration of a time-varying schedule. This definition allows us to sample $\xi$ independently from a continuous range during training while maintaining a consistent discrete schedule during the sampling phase.
@@ 3.2 Noise Conditional Score Networks
Typically, the gradients $\nabla_{\phi}\log p_{t}(\phi)$ in Eq. (9) are unknown. However, one can model a score function, $s(\phi,\xi)$, to match the gradients Song et al. (2020). Due to demands of the scalability and ability to efficiently handle high-dimensional parameter spaces, we employ a deep neural network to model the score function. To further improve sampling efficiency, we introduce a noise scheme for the forward process, in which $\xi_{i}$ ($i=1,\ldots,N_{T}$) is a positive geometric sequence that satisfies $\xi_{1}/\xi_{2}=\cdots=\xi_{N_{T}-1}/\xi_{N_{T}}>1$.
@@
$\forall\xi\in\{\xi_{i}\}_{i=1}^{N_{T}}:\quad s_{\theta}(\phi,\xi)\approx\nabla_{\phi_{\xi}}\log p_{\xi}(\phi_{\xi}|\phi_{0}),$ (10)
@@
$\ell(\theta;\xi)\triangleq\frac{1}{2}\mathbb{E}_{p_{\text{data}}(\phi)}\mathbb{E}_{\tilde{\phi}\sim\mathcal{N}(\phi,\sigma_{\xi}^{2}I)}\left[\left\lVert s_{\theta}(\tilde{\phi},\xi)+\frac{\tilde{\phi}-\phi_{0}}{\sigma_{\xi}^{2}}\right\rVert_{2}^{2}\right].$ (11)
@@
where $\lambda(\xi)=\sigma_{\xi}^{2}$ Song et al. (2020). Under this choice, we have
$\lambda(\xi)\ell(\theta;\xi)=\frac{1}{2}\mathbb{E}\left[\left\lVert\sigma_{\xi}s_{\theta}(\tilde{\phi},\xi)+\frac{\tilde{\phi}-\phi_{0}}{\sigma_{\xi}}\right\rVert_{2}^{2}\right].$ (13)
Since $(\tilde{\phi}-\phi_{0})/\sigma_{\xi}\sim\mathcal{N}(0,I)$ and empirically Song and Ermon (2019) $\lVert\sigma_{\xi}s_{\theta}(\phi,\xi)\rVert_{2}\propto 1$, we can deduce that the order of magnitude of $\lambda(\xi)\ell(\theta;\xi)$ is independent of $\xi$.
With regard to the specific network architecture, similar as in Ref. Wang et al. (2024), we adopt a fully-convolutional architecture with skip connection as shown in Fig. 3 to model the score function. It employs an encoder-decoder structure with symmetric skip connections. The receptive field grows as the network deepens, but each convolution acts locally, making it adaptable to arbitrarily large input sizes.
Figure 3: Net architecture. The gray lines with arrows denote skip connections between the encoding and decoding paths. In this study, the default input shape is specified as $(\cdot,2,16,16)$, where the first dimension represents the batch size, the second corresponds to the channel dimension, and the last two denote the spatial dimensions of the field configuration.
Input and Output. The input tensor is defined as $X\in\mathbb{R}^{N\times 2\times L\times L},$ where $N$ denotes the batch size, $2$ represents the two gauge link channels: angles of link variables in two dimensions, and $L$ (default $16$) is the spatial lattice size. The output tensor $Y$ has the same shape and approximates the score function, $s_{\theta}(\phi,\xi).$ For sampling, the lattice size $L$ can be adjusted accordingly.
Fully-Convolutional Net Architecture. The encoder applies convolutional layers,
$h^{(l+1)}=\sigma\left(W^{(l)}*h^{(l)}+b^{(l)}+d^{(l)}(e(\xi))\right),$ (14)
where the noise (time) dependence is introduced via an embedding function of the noise parameter $\xi$. Specifically, we first generate a Gaussian random feature embedding of $\xi$,
$e(\xi)=[\sin(2\pi W\xi),\cos(2\pi W\xi)],$ (15)
where $W$ in Eq. (3.9) is a shared “frequency embedding” matrix of shape
$W\in\mathbb{R}^{d_{\mathrm{embed}}\times d_{\xi}},$
where $d_{\xi}$ is the dimension of the noise/time input (typically $1$) and $d_{\mathrm{embed}}$ is the embedding dimension. This $W$ is shared across all layers and spatial locations (i.e. it does not vary per lattice site).
The function $d^{(l)}$ is implemented as a small fully-connected network,
$d^{(l)}:\mathbb{R}^{d_{\mathrm{embed}}}\to\mathbb{R}^{C_{\mathrm{out}}},$
where $C_{\mathrm{out}}$ is the number of feature channels. Its output is then broadcast spatially (i.e. replicated across all lattice sites) and added to the convolutional output at each spatial position. Thus, noise dependence explicitly modulates intermediate feature representations at each stage of the fully convolutional network.
Each layer follows group normalization, $\hat{h}^{(l)}={(h^{(l)}-\mu_{G})}/{\sqrt{\sigma_{G}^{2}+\epsilon}}.$, where $h^{(l)}$ denotes the input features at layer $l$, $\mu_{G}$ and $\sigma_{G}$ are respectively the mean and standard deviation computed over the feature channels for each normalization group $G$, and $\epsilon$ is a small constant added for numerical stability. The decoder utilizes transposed convolutions,
@@
where $E^{(l-1)}$ indicates the corresponding encoder layer. Finally, the network output is normalized by the standard deviation of the corresponding Gaussian noise.
Locality and Periodic Padding. The lattice gauge action is constructed from Wilson loops, ensuring local interactions. Accordingly, the fully convolutional network (FCN) computes outputs locally, as
@@
It denotes the collection (or local patch) of input values that lie within a neighbourhood $\mathcal{N}(i)$ around index $i$ and $\mathcal{N}(j)$ around index $j$. Periodic boundary conditions are enforced by using periodic padding, meaning that if an index exceeds the lattice boundaries, it wraps around according to the modulo operation. We note that due to the intrinsic periodic boundary conditions of the lattice, our setup differs from the conventional U-Net used in diffusion models. In contrast to the standard U-Net, which progressively reduces and then recovers spatial resolution to combine features of different granularities via skip connections, we keep the lattice size fixed throughout the network. This design naturally respects the periodic structure of the gauge fields.